Please help I am not sure how to solve the following!
s=1/2 gt^2 (solve for g)

Found 2 solutions by longjonsilver, Scorinitron:
Answer by longjonsilver(2297)

look at my percentage Lesson on this site

Jon.

Answer by Scorinitron(19)

s=(1)/(2)*gt^(2)

Since g is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
(1)/(2)*gt^(2)=s

Multiply (1)/(2) by gt^(2) to get (gt^(2))/(2).
(gt^(2))/(2)=s

Multiply each term in the equation by 2.
(gt^(2))/(2)*2=s*2

Simplify the left-hand side of the equation by canceling the common factors.
gt^(2)=s*2

Multiply s by 2 to get 2s.
gt^(2)=2s

Divide each term in the equation by t^(2).
(gt^(2))/(t^(2))=(2s)/(t^(2))

Cancel the common factor of t^(2) in (gt^(2))/(t^(2)).

g=(2s)/(t^(2))